OpenStudy (anonymous):

what is the equation of the line that passes through the point (2,5) and has a slope of 2? a. y=2x+1 b.y=2x+3 c.y=2x-8 d.y=2x-3

OpenStudy (anonymous):

answer : a (x,y) = (2,5) slope = 2 = gradient, m y = mx + c 5 = 2 (2) + c c = 1 substitute back c = 1 and m = 2, eq : y = 2x + 1

OpenStudy (anonymous):

The answer is 'a'